Identity, Self-righteousness and COPS

Matthew Gray

April 15, 1997

Introduction

The reasons for the success of the COPS television show are not entirely clear. FOX Television attributes its success and appeal to the fact that it is “unpredictable”, “riveting”, and “an intimate, up-close look at life as seen through the eyes of the brave police officers who patrol our cities.” While these claims are questionable (certainly the “unpredictable” claim), they don’t provide any deep understanding of the shows appeal.

Seemingly obvious reasons for the shows popularity don’t hold up terribly well. One might claim the shows “excitement” attracts viewers. Much of the time, however, COPS is not at all exciting, and the segments revolve around minor, often mundane criminal activity. Another possibility is a certain civil appreciation for law enforcement, however after viewing the show a few times, this seems rather unlikely.

Identity

One possible, less obvious, alternative explanation for the show’s success is a sense of identification (and anti-identification) people have with the show. Some viewers who experience many of the minor to moderate crimes such as those portrayed on COPS on a regular basis, can identify with the police. The police are thrust into a more extreme version of the viewers’ lives, complete with mundane, but unpleasant crime. The police, unlike the viewers, are capable of taking quick, decisive legal action against the perpetrators, and in COPS , reliably succeed.

Equally so, there is, in some cases a sense of identity with the perpetrators, though much less so. Particularly in cases where the police in the show are harassing or pursuing a suspect who is not “obviously” guilty, viewers may identify with the suspect. This provides an easy model for the viewer to cast themselves into the role of the suspect, and the police into the role of any authority figure or even society at large. Seeing others in what is perceived as a similar plight enhances the perception that theirs is a case of injustice rather than justice they have misconstrued.

Finally, there is a certain amount of “anti-identification” with both the criminals and the the police, though more so the criminals. This manifests itself as an outrage at the behavior and a strong sense of identity that the viewer is not like the people portrayed.

Self-righteousness

Related to, but different from this sense of anti-identification is a sense of self-righteousness. This occurs in viewing both the police and the suspects. In the case of the suspects, this primarily takes form as thoughts to the effect of “I’m smarter than that.” or “I’m a better person than them because I’m not a criminal”. Watching suspects “get busted” reaffirms feelings that the fact that they are a law-abiding citizen makes them “better” in some way.

Similarly, there is a sense when viewing the police of “I could do that.” or even, “I could do that better”. Though the status of police has changed for the worse in recent decades, it is still a relatively respeced and prestigious position, particularly for its required level of education. Viewers perception that they are equals or better to police in competence, intelligence, or capability enhances the feelings of self-righteousness.

Overall, self-righteousness as an emotion helps the viewer identify their place in society, and affirm them that it is sufficiently high. Very much, the over all effect is the opposite of many other shows, which enhance people’s lives by providing vicarious experience. Instead, COPS provides an experience that is almost familiar and enhances their sense of identity without diverging too far from their actual lives.

Conclusion

The mundane but familiar qualities of COPS have contributed to its success. The fact that it is “real” further cements its role as a concrete societal benchmark against which viewers measure themselves. This is much of what happens with normal television, however COPS succeeds in being a compelling benchmark more than most. It does so by being realistic and not setting the mark too high, and keeping a fairly clear distinction in the two measures it provides. The criminals are reliably “wrong” and often stupid. The police are successful, but rearely outrageously so, and are reliably conspicously imperfect. This allows essentially all viewers to measure themselves above the criminals, and often above the police, but usually on par with them.